OpenStudy (anonymous):

surface area when y = 1/4 X^2 - 1/2 lnx is revolved about the y axis.

OpenStudy (jamesj):

between what values of x or y?

OpenStudy (anonymous):

1<x<2

OpenStudy (jamesj):

Ok, so in general for a function y = f(x) what is the surface area of rotation for x between a and b? What is the integral we want to evaluate?

OpenStudy (anonymous):

i got to ds = 2\[2\pi \int\limits_{1}^{2}x \sqrt{1 + (x - 1/x)^{2}}\]

OpenStudy (jamesj):

Now are you rotating about the x axis or y axis?

OpenStudy (anonymous):

y axis